OpenStudy (anonymous):

Factor 3x^2 - 9x - 20 Factor 3x^2 - 9x - 20 @Mathematics

OpenStudy (asnaseer):

this cannot be factored - are you sure you have the correct equation?

OpenStudy (asnaseer):

the closest I can get are:\[(3x+5)(x-4)=3x^{2}-7x-20\]or\[(3x+4)(x-5)=3x^{2}-11x-20\]

OpenStudy (anonymous):

Yeah I have to find x from this equation: (2 / x - 2) + (7 / x^2 - 4) = (5/x)

OpenStudy (asnaseer):

then you need to use the quadratic equation solution to find the two roots of x

OpenStudy (anonymous):

I can't use the quadratic equation on the that, the equation i gave is a rational equation

OpenStudy (asnaseer):

I just worked out the equation from the new information you gave and it should be:\[3x^{2}-11x-20=0\]

OpenStudy (asnaseer):

so solution is:\[(3x+4)(x-5)=0\]which gives x=5 or -4/3

OpenStudy (anonymous):

OHH! I got it lol i realized I didnt distribute correctly lol